OpenStudy (anonymous):

I just do not understand some teachers....take a look at this and tel me what you think...

OpenStudy (anonymous):

My sis is in 7th grade. She got a paper with problems on it and I checked them and to me they seemed to be correct....Here is one of them...and this is how my sis answered it... (12 divided by 3) + (16 divided by 4) 4 + 4 = 8 My sis turned it in and the teacher said she would have to stay for detention because she didn't follow instructions by working the problem out. Anything that has to be worked out in this problem can be done in your head. I do not get it.

OpenStudy (anonymous):

should she have used long division and worked out 12 divided by 3 and 16 divided by 4 ??

OpenStudy (anonymous):

Some teachers want to see every step. Your sister did not show the translation between words and math terms. She simplified directly because it could be done in her head. Perhaps she should have written out 12/3+14/4= 4+4=8???

OpenStudy (anonymous):

I guess so.....picky....she got the right answer....now she is having to stay for detention.

OpenStudy (anonymous):

I told her to ask the teacher exactly how she was supposed to work out the problems. I am curious as to what the teacher is gonna say

OpenStudy (anonymous):

That's why some kids hate mathematics. Tell your sister to hang in there.

OpenStudy (anonymous):

thanks...I will